# Large Deviation Principles for Abelian Monoids

### v1

@@ Abstract.
Following work of Mehrdad and Zhu [MZ16] and of Liu [Liu04a, Liu04b], we prove a large deviation principle for a broad class of integer-valued additive functions defined over abelian monoids. As a corollary, we obtain a large deviation principle for a generalized form of the Erdös-Kac theorem due to Liu.
@@ 1. Introduction
The celebrated Erdős-Kac Theorem [EK40] says that if $\omega(n)$ is the number of distinct prime factors of an integer $n$, then
@@
Much work has been done on the rate at which (1) converges to $N(0,1)$, see [RT58] for one such example. Likewise, one can study the tails of the distribution as the upper bound for $n$ grows. Indeed, in [MZ16], Mehrdad and Zhu prove a large deviation principle for a large class of “strongly additive” functions $g:\mathbb{N}\to\mathbb{N}$, including when $g$ is the number of prime divisors function, $\omega$, thus giving a large deviation principle for the distributions of (1) over $n\leq X$ as $X\to\infty$.
@@
The application we have in mind is when $g$ is the number-of-distinct-prime-divisors function, $\omega$. For $m=\prod_{p\in\mathcal{P}}p^{e_{p}}\in\mathcal{M}$, let $\omega(m)$ be the number of $p\in\mathcal{P}$ for which $e_{p}\geq 1$.
@@
For each $X\in\mathbb{N}$, we denote by $V(X)$ a uniformly chosen monoid element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$. For each $p\in\mathcal{P}$, let $Z$ be a random variable such that $Z_{p}=1$ if $V(X)$ is divisible by $p$, and $Z_{p}=0$ otherwise.
@@
where $V(n)$ is a uniformly chosen monoidal element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$, see Definition 1.4. Then for any Borel measurable set $A\subset\mathbb{R}$, the probability $\mathbb{P}\left[\frac{W(X)}{\log\log X}\in A\right]$ satisfies a large deviation principle with speed $\log\log X$ and rate function
@@ 3. Preliminary Results
$-\inf_{x\in A^{o}}I(x)\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]$
@@
For distinct $p_{1},...,p_{k}\in\mathcal{P}$ we have
@@
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\sum_{m\in\mathcal{M}}\mathbbm{1}_{p_{1}p_{2}\cdots p_{k}\mid m}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}}$
@@
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]=\frac{1}{aX}\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O\left(\max\left(\frac{aX^{b-1}}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})},X^{2b-2}\right)\right).$ (8)
@@
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]\leq\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}]+O\left(\max\left(\frac{aX^{b-1}}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})},X^{2b-2}\right)\right).$
@@
    For any non-negative sequence of real numbers $\{\theta_{p}\}_{p\in\mathcal{P}}$ we have
    $\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Z_{p}\right)\right]\leq\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Y_{p}+O\left(\max\left(\frac{aX^{b-1}}{\prod_{\mathbf{N}(p)\leq X}\mathbf{N}(p)}\right),X^{2b-2}\right)\right)\right].$
All the implied constants of the error terms are independent of the choice of $p_{1},\cdots,p_{k}$.
@@
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O(X^{b})}{aX+O(X^{b})}$
$=\frac{1}{aX}\left\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\right\rfloor+O\left(\max\left(\frac{aX^{b-1}}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})},X^{2b-2}\right)\right).$
@@
The third statement follows by using Taylor expansion. ∎
@@
By the exponential [^1] For an integrable random variable $X$ with finite, non-zero variance and $\varepsilon>0$, $P(X\geq\varepsilon)\leq e^{-\theta\varepsilon}\mathbb{E}(e^{\theta X})$ for any $\theta>0.$ Chebyshev’s Inequality with any $\theta>0$, and Lemma 3.4, we have
$\begin{split}&\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\geq\varepsilon\log\log X\right)\\ &\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{E}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\right)\right]-\theta\varepsilon\\ &\leq\limsup_{X\to\infty}\log\mathbb{E}\frac{1}{\log\log X}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]-\theta\varepsilon\end{split}$ (11)
[…]
@@
$\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\leq\epsilon\log\log X\right]=-\infty.$ (14)
@@
By Condition 1.1 and Lemma 3.4, for any $\theta>0$ we have
$\hskip 14.0pt\log\mathbb{E}\left[\exp\left(\theta\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\right)\right]$
$\leq\log\mathbb{E}\left[\exp\left(\theta C\sum_{p\in A(X,C)}Y_{p}\right)\right]+O\left(\max\left(\frac{aX^{b-1}}{\prod_{\mathbf{N}(p)\leq X}\mathbf{N}(p)},X^{2b-2}\right)\right)$
$=\sum_{p\in A(X,C)}\log\left((\exp\left(\theta C\right)-1)\cdot\frac{1}{\mathbf{N}(p)}+1\right)+O\left(\max\left(\frac{aX^{b-1}}{\prod_{\mathbf{N}(p)\leq X}\mathbf{N}(p)},X^{2b-2}\right)\right)$
$\leq(\exp(\theta C)-1)\cdot\sum_{k_{X}\leq\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}+O\left(\max\left(\frac{aX^{b-1}}{\prod_{\mathbf{N}(p)\leq X}\mathbf{N}(p)},X^{2b-2}\right)\right).$
@@
$\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\leq\epsilon\log\log X\right]$
@@
We let $\theta$ to grow arbitrarily large to conclude the lemma. ∎
@@
By Lemma 3.4, for any tuples of non-negative integers $r_{1},\cdots,r_{k}$, there exists an absolute constant $B>0$ such that
@@
Note that we have
(17)(18)$+O\left(\max\left(\frac{aX^{b-1}}{\prod_{p\in B(X,C)}\mathbf{N}(p)},X^{2b-2}\right)\right),$ ) ) ,
@@
$\leq\lim_{X\to\infty}B^{\prime\prime}e^{-\frac{1}{2}(1-b)\log X}+2B_{**}e^{-\frac{1}{2}\log K\log\log X}=0.$
@@ 5. Acknowledgments
This work was completed while the first named author visited the second at the Max Planck Institute for Mathematics in Bonn; we thank MPIM for its generous support and for providing excellent working conditions.

### v2

@@ Abstract.
Following work of Mehrdad and Zhu [MZ16] and of Liu [Liu04a, Liu04b], we prove a large deviation principle for a broad class of integer-valued additive functions defined over abelian monoids. As a corollary, we obtain a large deviation principle for a generalized form of the Erdős-Kac theorem due to Liu.
@@ 1. Introduction
The celebrated Erdős-Kac Theorem [EK40] says that if $\omega(n)$ is the number of distinct prime factors of a positive integer $n$, then
@@
Much work has been done on the rate at which (1) converges to $N(0,1)$, see [RT58] for one such example. Likewise, one can study the tails of the distribution as the upper bound for $n$ grows. Indeed, in [MZ16], Mehrdad and Zhu prove a large deviation principle for a large class of “strongly additive” functions $g:\mathbb{N}\to\mathbb{N}$, including when $g$ is the number-of-distinct-prime-divisors function, $\omega$, thus giving a large deviation principle for the distributions of (1) over $n\leq X$ as $X\to\infty$.
@@
The application we have in mind is when $g$ is the number-of-distinct-prime-divisors function, $\omega$. On $\mathcal{M}$, this means for $m=\prod_{p\in\mathcal{P}}p^{e_{p}}$, let $\omega(m)$ be the number of $p\in\mathcal{P}$ for which $e_{p}\geq 1$.
@@
For each $n\in\mathbb{N}$, we denote by $V(n)$ a uniformly chosen monoid element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq n\}$. For each $p\in\mathcal{P}$, let $Z$ be a random variable such that $Z_{p}=1$ if $V(n)$ is divisible by $p$, and $Z_{p}=0$ otherwise.
@@
where $V(n)$ is a uniformly chosen monoidal element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq n\}$, see Definition 1.4. Then for any Borel measurable set $A\subset\mathbb{R}$, the probability $\mathbb{P}\left[\frac{W(X)}{\log\log X}\in A\right]$ satisfies a large deviation principle with speed $\log\log X$ and rate function
@@ 3. Preliminary Results
$-\inf_{x\in A^{o}}I(x)\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]\leq-\inf_{x\in\overline{A}}I(x)$
@@
Given $p\in\mathcal{P}$, we recall the random variable $Z_{p}$ supported over the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$ from Definition 1.4. For distinct $p_{1},...,p_{k}\in\mathcal{P}$ we have
@@
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\sum_{\begin{subarray}{c}m\in\mathcal{M}\\ \mathbf{N}(m)\leq X\end{subarray}}\mathbbm{1}_{p_{1}p_{2}\cdots p_{k}\mid m}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}}$
@@
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]=\frac{1}{aX}\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
@@
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]\leq\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}]+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
@@
    For sufficiently large $X$ there exists an explicit positive constant $M$, independent of the choice of distinct primes $p_{1},\cdots,p_{k}\in\mathcal{P}$, such that
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]\leq M\cdot\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}].$
  4. (4)
    For sufficiently large $X$, there exists an explicit constant $M$ such that for any non-negative sequence of real numbers $\{\theta_{p}\}_{p\in\mathcal{P}}$ we have
    $\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Z_{p}\right)\right]\leq M\cdot\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Y_{p}\right)\right].$
The implied constants of all the error terms are independent of the choice of $p_{1},\cdots,p_{k}$.
@@
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O\left(\frac{X^{b}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right)}{aX+O(X^{b})}$
$=\frac{1}{aX}\left\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\right\rfloor+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
@@
To prove the third statement, we divide into two cases. Suppose $p_{1},p_{2},\cdots,p_{k}$ are distinct primes such that $\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})>X$. Then because all the random variables $Z_{p_{1}},Z_{p_{2}},\cdots,Z_{p_{k}}$ are supported over the probability space $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$, we have
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]=0<\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}=\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
Now suppose $p_{1},p_{2},\cdots,p_{k}$ are distinct primes such that $\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})\leq X$. Then we have
$\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}$ $=\frac{1}{X^{1-b}\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}$
$\leq\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}=\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
Hence, by the second statement of our lemma, there exists an absolute constant $M_{1}>0$ such that
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]\leq M_{1}\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
We take $M:=\max(1,M_{1})$ to obtain the third statement of the lemma.
The fourth statement follows by using Taylor expansion. We crucially use the fact that for any non-negative integers $r_{1},r_{2}\cdots,r_{k}$ we have
$\mathbb{E}[Z_{p_{1}}^{r_{1}}Z_{p_{2}}^{r_{2}}\cdots Z_{p_{k}}^{r_{k}}]=\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]\leq M\cdot\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}]=M\cdot\mathbb{E}[Y_{p_{1}}^{r_{1}}Y_{p_{2}}^{r_{2}}\cdots Y_{p_{k}}^{r_{k}}].$
@@
By the exponential [^1] For an integrable random variable $X$ with finite, non-zero variance and $\varepsilon>0$, $P(X\geq\varepsilon)\leq e^{-\theta\varepsilon}\mathbb{E}(e^{\theta X})$ for any $\theta>0.$ Chebyshev’s Inequality with any $\theta>0$, and the third statement of Lemma 3.4, there exists an explicit constant $M>0$ such that
$\begin{split}&\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\geq\varepsilon\log\log X\right)\\ &\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{E}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\right)\right]-\theta\varepsilon\\ &\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\left(\log\mathbb{E}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]+\log M\right)-\theta\varepsilon\\ &=\limsup_{X\to\infty}\log\mathbb{E}\frac{1}{\log\log X}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]-\theta\varepsilon.\end{split}$ (10)
@@
$\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\geq\epsilon\log\log X\right]=-\infty.$ (13)
@@
By Condition 1.1 and the third statement of Lemma 3.4, for any $\theta>0$ and sufficiently large $X$ there exists an explicit constant $M>0$ such that
$\log\mathbb{E}\left[\exp\left(\theta\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\right)\right]$ $\leq\log\mathbb{E}\left[\exp\left(\theta C\sum_{p\in A(X,C)}Y_{p}\right)\right]+\log M$
$=\sum_{p\in A(X,C)}\log\left((\exp\left(\theta C\right)-1)\cdot\frac{1}{\mathbf{N}(p)}+1\right)+\log M$
$\leq(\exp(\theta C)-1)\cdot\sum_{k_{X}\leq\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}+\log M.$
@@
$\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\geq\epsilon\log\log X\right]$
@@
Let $\theta$ grow arbitrarily large to conclude the lemma. ∎
@@
By the second statement of Lemma 3.4, for any tuples of non-negative integers $r_{1},\cdots,r_{k}$ and sufficiently large $X$ there exists an absolute constant $B>0$ such that
@@
Note that the third statement of Lemma 3.4 implies that for sufficiently large $X$ there exists an explicit constant $M>0$ such that
[…]
@@
$\leq\lim_{X\to\infty}B^{\prime\prime}\cdot e^{-\frac{1}{2}(1-b)\log X}+(M+1)\cdot B_{**}\cdot e^{-\frac{1}{2}\log K\log\log X}=0.$
@@ 5. Acknowledgments
This work was completed while the first author visited the second at the Max Planck Institute for Mathematics in Bonn; we thank MPIM for its generous support and for providing excellent working conditions. We also thank Sachin Kumar for pointing out a typo in the main theorem.